nolegirl24:

Ik the first part but i forgot how to do the answer choices What is the exact value for the expression the square root of 52. − the square root of 13. + the square root of 117.? Simplify if possible. 4 the square root of 13. the square root of 13. 2 the square root of 39. 8 the square root of 39.

Vocaloid:

so you want to re-write each product as (something)*13

Vocaloid:

52 = __ * 13 fill in the blank

nolegirl24:

2.17????? Do you do 13/52

Vocaloid:

other way around

nolegirl24:

Oh okay hold on

Vocaloid:

you would divide both sides by 13 to get 52/13 = ?

nolegirl24:

4

Vocaloid:

good so sqrt(52) = sqrt(4*13) = 2sqrt(13)

Vocaloid:

now let's do the same thing with 117 117 = __ * 3

nolegirl24:

39

Vocaloid:

oh whoops made a typo

Vocaloid:

117 = ___ * 13

Vocaloid:

try that ;;

nolegirl24:

39

Vocaloid:

117/13 = ?

nolegirl24:

Sorry 9 not 39

Vocaloid:

good so sqrt(117) = sqrt(9*13) = 3(sqrt13) so now they're all in terms of sqrt(13) we just add/subtract like terms

Vocaloid:

square root of 52. − the square root of 13. + the square root of 117. 2sqrt(13) - sqrt(13) + 3sqrt(13) = 4(sqrt13) = answer A

nolegirl24:

Thank you!